OpenStudy (anonymous):

Find what angle MPQ is, if the arc is 9 and the radius is 5

OpenStudy (anonymous):

If the arc length is 9 and the radius is 5, then --> pi*5*2*x = 9 x = .9*pi degrees

OpenStudy (anonymous):

that makes no sence

OpenStudy (anonymous):

That really doesn't make any sense.

OpenStudy (anonymous):

Sorry I messed it up hang on.

OpenStudy (anonymous):

Okay so let g be the ratio of the arc to the entire circle. a is the angle, so g = a/360 Then, \[9 = 2\pi*r*g\] \[9 = 2*\pi*r*a/360\] \[a = 9/(10*\pi) * 360\] = 103 degrees

OpenStudy (anonymous):

thank you